# frame=single, fontsize=, breaklines

### v2

@@ 4.1 Python Library
Users can install the library by simply using the pip package manager: {minted}bash $pipinstallopenfactcheck\par Thelibraryincludesdetaileddocumentationtoassistdevelopersincustomizingandextendingthefunctionalitytomeettheirspecificneedsanditiscontinuallyupdatedtoensurecompatibilitywiththelatestresearchanddatasecuritystandards.\par$
@@
The first step is to import the necessary library components and initialize OpenFactCheckConfig configuration and OpenFactCheck class, which requires no input values for default usage, as shown below: {minted}python from openfactcheck import OpenFactCheck, OpenFactCheckConfig config = OpenFactCheckConfig() ofc = OpenFactCheck(config)
@@
{minted}
python ofc.ResponseEvaluator.evaluate(response: str) # response: string output from LLM
{minted}
python ofc.LLMEvaluator.evaluate(model_name: str, input_path: str) # model_name: evaluated model name. # input_path: path to the CSV containing responses for the LLM Benchmark.
# Output # A dictionary with detailed scores (precision, recall, f1, accuracy, cost, time etc. for each dataset subset i.e. snowballing, selfaware, freshqa, factoolqa, felm-wk, factcheck-bench and factscore-bio.
{minted}
python ofc.CheckerEvaluator.evaluate(input_path: str) # input_path: path to the CSV containing responses for the FactChecker Benchmark
# Output # A dictionary with detailed scores (precision, recall, f1, accuracy, cost, time etc.)
@@ Appendix A Pseudo Code of ResponseEval
{minted}
@@
# FacTool claims = call_LLM(document, promot="extract context-independent atomic claims based on the document")
return claims
def retriever(claim: str, database: DB, retrieval_strategy: obj, search_api_key: str) -> List[str]: # offline DB dump evidence = retrieval_strategy(claim, database)
# online web pages by calling API evidence = serper_or_serpapi(claim, search_api_key)
return evidence
def verifier(claim: str, evidence: List[str]) -> bool: # call LLMs factual_label = call_LLM(claim, evidence, prompt="based on the evidence and your own knowledge, determine whether the claim is true or false.")
# use NLI models stance2factual = "entailment": true, "contradiction": false, "neutral": "not enough evidence" stances = [nli(evid, claim) for evid in evidence] majority_stance = majority_vote(factual_labels) factual_label = stance2factual[majority_stance]
return factual_label

### v3

# OpenFactCheck: A Unified Framework for Factuality Evaluation of LLMs

@@ 4.1 Python Library
Users can install the library by simply using the pip package manager:
\$ pip install openfactcheck
The library includes detailed documentation to assist developers in customizing and extending the functionality to meet their specific needs and it is continually updated to ensure compatibility with the latest research and data security standards.
@@
The first step is to import the necessary library components and initialize OpenFactCheckConfig configuration and OpenFactCheck class, which requires no input values for default usage, as shown below:
from openfactcheck import OpenFactCheck, OpenFactCheckConfig
config = OpenFactCheckConfig()
ofc = OpenFactCheck(config)
@@
ofc.ResponseEvaluator.evaluate(response: str)
# response: string output from LLM
ofc.LLMEvaluator.evaluate(model_name: str, input_path: str)
# model_name: evaluated model name.
# input_path: path to the CSV containing responses for the LLM Benchmark.
[…]
# A dictionary with detailed scores (precision, recall, f1, accuracy, cost, time etc. for each dataset subset i.e. snowballing, selfaware, freshqa, factoolqa, felm-wk, factcheck-bench and factscore-bio.
ofc.CheckerEvaluator.evaluate(input_path: str)
# input_path: path to the CSV containing responses for the FactChecker Benchmark
[…]
# A dictionary with detailed scores (precision, recall, f1, accuracy, cost, time etc.)
@@ Appendix A Pseudo Code of ResponseEval
def claim_processor(document: str) -> List[str]:
# FactScore
@@
sentences = [NLTK(para) for para in paragraphs]
claims = [call_LLM(sentence, prompt="decompose into atomic claims") for sentence in sentences]
[…]
claims = call_LLM(document, promot="extract context-independent atomic claims based on the document")
[…]
# offline DB dump
evidence = retrieval_strategy(claim, database)
[…]
evidence = serper_or_serpapi(claim, search_api_key)
[…]
# call LLMs
factual_label = call_LLM(claim, evidence, prompt="based on the evidence and your own knowledge, determine whether the claim is true or false.")
[…]
stance2factual = {
"entailment": true,
"contradiction": false,
"neutral": "not enough evidence"
stances = [nli(evid, claim) for evid in evidence]
majority_stance = majority_vote(factual_labels)
factual_label = stance2factual[majority_stance]
[…]